OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

OpenStudy (anonymous):

you want the distance ?

OpenStudy (anonymous):

here are my options: y = 2x – 8 y = – 2x + 9 y = 2x – 11 y = – 2x + 6

OpenStudy (anonymous):